# On Jiang’s wavefront sets conjecture for representations in local Arthur packets

### 13 Mar 2026 v1

@@ 5. Construction of elements in local Arthur packets
Then we know that
$\dim_{\mathfrak{g}_{n}}(\underline{p})>\dim_{\mathfrak{g}_{n}}(\eta_{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}(\underline{p}(\psi))).$
From identities (6.1) and (6.2), we know that for any nilpotent orbit $\mathcal{O}$},
$c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})=\sum_{\pi\in\widetilde{\Pi}_{\psi}}\langle 1,\pi\rangle\omega_{\pi}(c_{{\mathrm{G}}_{n}})c_{{\mathcal{O}}}(\pi)=\sum_{\pi\in\widetilde{\Pi}_{\psi}}\omega_{\pi}(c_{{\mathrm{G}}_{n}})c_{{\mathcal{O}}}(\pi).$
On the other hand, since $\underline{p}\in max\{\cup_{\pi\in\widetilde{\Pi}_{\psi}}\mathfrak{p}^{m}(\pi)\}$, by [MW87], for any nilpotent orbit ${\mathcal{O}}_{\underline{p}}$ corresponding to $\underline{p}$, $c_{{\mathcal{O}}_{\underline{p}}}(\pi)\in\mathbb{Z}_{\geq 0}$ and there exist $\pi\in\widetilde{\Pi}_{\psi}$ and a rational nilpotent orbit ${\mathcal{O}}_{0}\in{\mathcal{N}}_{\mathfrak{g}_{n}}$ corresponding to $\underline{p}$ such that $c_{{\mathcal{O}}_{0}}(\pi)\in\mathbb{Z}_{>0}$. By Proposition A.4, all representations in the local Arthur packets $\widetilde{\Pi}_{\psi}$ have the same central character. Therefore, $c_{{\mathcal{O}}_{0}}(\widetilde{\Pi}_{\psi})\neq 0$. By Lemma 6.1, there exists a function $f_{0}\in\mathcal{H}(G_{n})$ with arbitrarily small support such that
@@
for any other ${\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}$. Let $f=R(c_{{\mathrm{G}}_{n}})f_{0}$ when $\mathrm{G}_{n}={\mathrm{Sp}}_{2n},{\mathrm{SO}}_{2n+1}$, and let $f=R(c_{{\mathrm{G}}_{n}})(f_{0}+f_{0}^{c})$ when $\mathrm{G}_{n}={\mathrm{SO}}_{2n}^{\alpha}$, where $c$ is the non-trivial element in the outer automorphism group. Then $f\in\widetilde{{\mathcal{H}}}(G_{n})$. When $\mathrm{G}_{n}={\mathrm{SO}}_{2n}^{\alpha}$, if ${\mathcal{O}}_{0}\neq{\mathcal{O}}_{0}^{c}$, i.e., ${\mathcal{O}}_{0}$ is a very even nilpotent orbit, then we have that
$\hat{\mu}_{{\mathcal{O}}_{0}}(R(c_{{\mathrm{G}}_{n}})f)=\hat{\mu}_{{\mathcal{O}}_{0}^{c}}(R(c_{{\mathrm{G}}_{n}})f)=1,\text{ and }\hat{\mu}_{{\mathcal{O}}}(R(c_{{\mathrm{G}}_{n}})f)=0,$
@@
Now, without loss of generality, we can assume that $R(c_{{\mathrm{G}}_{n}})f$, $R(c_{{\mathrm{G}}_{1}^{\prime}})f^{\prime 1}\otimes R(c_{{\mathrm{G}}_{2}^{\prime}})f^{\prime 2}$, $R(s_{\widehat{\mathrm{G}}_{1}^{\prime}})\widetilde{f}^{1}$, $R(s_{\widehat{\mathrm{G}}_{2}^{\prime}})\widetilde{f}^{2}$, $\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}$, and $\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}$ all have small supports near 1. Take $t\in F^{\times}$ with $|t|$ small. By Lemma 6.2 and Lemma 6.3, when ${\mathrm{G}}_{n}={\mathrm{Sp}}_{2n}$, we have
(6.3) $\begin{split}&\,\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{\mathcal{O}}((R(c_{{\mathrm{G}}_{n}})f)_{t^{2}})\\ =&\,|t|^{\dim(\mathfrak{g}_{n})-\dim(\mathfrak{g^{\prime}})}\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{{\mathfrak{g}}_{1}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}((\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}})_{t^{2}})\right)\\ &\,\cdot\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}((\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}})_{t^{2}})\right);\end{split}$
when ${\mathrm{G}}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}$, we have
(6.4) $\begin{split}&\,\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{\mathcal{O}}((R(c_{{\mathrm{G}}_{n}})f)_{t^{2}})\\ =&\,|t|^{\dim(\mathfrak{g}_{n})-\dim(\mathfrak{g^{\prime}})}\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{1}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}((\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}})_{t^{2}})\right)\\ &\,\cdot\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}((\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}})_{t^{2}})\right).\end{split}$
@@
$\hat{\mu}_{\mathcal{O}}((R(c_{{\mathrm{G}}_{n}})f)_{t^{2}})$ $=|t|^{2\dim(\mathfrak{g}_{n})-\dim({\mathcal{O}})}\hat{\mu}_{\mathcal{O}}(R(c_{{\mathrm{G}}_{n}})f),$
$\hat{\mu}_{\mathcal{O}}((\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}})_{t^{2}})$ $=|t|^{2\dim(\hat{\mathfrak{g}}_{1}^{\prime})-\dim({\mathcal{O}})}\hat{\mu}_{\mathcal{O}}(\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}),$
$\hat{\mu}_{\mathcal{O}}((\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}})_{t^{2}})$ $=|t|^{2\dim(\hat{\mathfrak{g}}_{2}^{\prime})-\dim({\mathcal{O}})}\hat{\mu}_{\mathcal{O}}(\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}).$
Also, when ${\mathrm{G}}_{n}={\mathrm{Sp}}_{2n}$,
$\hat{\mu}_{\mathcal{O}}((\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}})_{t^{2}})=|t|^{2\dim({\mathfrak{g}}_{1}^{\prime})-\dim({\mathcal{O}})}\hat{\mu}_{\mathcal{O}}(\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}).$
Hence, from (6.3), we obtain that
@@
(6.5) $\begin{split}&\,|t|^{\dim(\mathfrak{g}_{n})-\dim({\mathcal{O}}_{0})}c_{{\mathcal{O}}_{0}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{{\mathcal{O}}_{0}}(R(c_{{\mathrm{G}}_{n}})f)\\ =&\,\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{{\mathfrak{g}}_{1}^{\prime}}}|t|^{\dim({\mathfrak{g}}_{1}^{\prime})-\dim({\mathcal{O}})}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}(\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}})\right)\\ &\,\cdot\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}}|t|^{\dim(\hat{\mathfrak{g}}_{2}^{\prime})-\dim({\mathcal{O}})}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}(\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}})\right);\end{split}$
from (6.4), we obtain that
(6.6) $\begin{split}&\,|t|^{\dim(\mathfrak{g}_{n})-\dim({\mathcal{O}}_{0})}c_{{\mathcal{O}}_{0}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{{\mathcal{O}}_{0}}(R(c_{{\mathrm{G}}_{n}})f)\\ =&\,\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{1}^{\prime}}}|t|^{\dim(\hat{\mathfrak{g}}_{1}^{\prime})-\dim({\mathcal{O}})}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}(\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}})\right)\\ &\,\cdot\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}}|t|^{\dim(\hat{\mathfrak{g}}_{2}^{\prime})-\dim({\mathcal{O}})}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}(\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}})\right).\end{split}$
@@
Then one can see that when ${\mathrm{G}}_{n}={\mathrm{Sp}}_{2n}$, there exist nilpotent orbits ${\mathcal{O}}_{1}\in{\mathcal{N}}_{{\mathfrak{g}}_{1}^{\prime}}$, ${\mathcal{O}}_{2}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}$ such that $c_{{\mathcal{O}}_{1}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{{\mathcal{O}}_{1}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})$ and $c_{{\mathcal{O}}_{2}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{{\mathcal{O}}_{2}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})$ are all nonzero and
$\begin{split}\dim(\mathfrak{g}_{n})-\dim_{\mathfrak{g}_{n}}({\mathcal{O}}_{0})=\dim({\mathfrak{g}}_{1}^{\prime})+\dim(\hat{\mathfrak{g}}_{2}^{\prime})-\dim_{{\mathfrak{g}}_{1}^{\prime}}({\mathcal{O}}_{1})-\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}({\mathcal{O}}_{2}).\end{split}$
When ${\mathrm{G}}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}$, there exist nilpotent orbits ${\mathcal{O}}_{i}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{i}^{\prime}}$ such that $c_{{\mathcal{O}}_{i}}(\widetilde{\pi}_{\psi^{i}})\hat{\mu}_{{\mathcal{O}}_{i}}({{}\widetilde{f}^{i}_{\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}}})\neq 0$, $i=1,2$, and
$\begin{split}\dim(\mathfrak{g}_{n})-\dim_{\mathfrak{g}_{n}}({\mathcal{O}}_{0})=\dim(\hat{\mathfrak{g}}_{1}^{\prime})+\dim(\hat{\mathfrak{g}}_{2}^{\prime})-\dim_{\hat{\mathfrak{g}}_{1}^{\prime}}({\mathcal{O}}_{1})-\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}({\mathcal{O}}_{2}).\end{split}$
When ${\mathrm{G}}_{n}={\mathrm{Sp}}_{2n}$, $\mathfrak{g}_{1}^{\prime}=\mathfrak{sp}_{2n_{1}}$, $\hat{\mathfrak{g}}_{1}^{\prime}=\mathfrak{so}_{2n_{1}+1}$, hence, $\dim(\mathfrak{g}_{1}^{\prime})=\dim(\hat{\mathfrak{g}}_{1}^{\prime})$. By Lemma 4.1 with $J=\emptyset$, or by Proposition 4.2,
$\dim_{{\mathfrak{g}}_{1}^{\prime}}(\eta_{{\hat{\mathfrak{g}}_{1}^{\prime},\mathfrak{g}_{1}^{\prime}}}(\underline{p}(\psi^{1})))=\dim_{\hat{\mathfrak{g}}_{1}^{\prime}}((\underline{p}(\psi_{1})^{t})_{\widehat{\mathrm{G}}_{1}^{\prime}}).$
By Lemma 4.1 again, we can see that
$\dim_{{\mathfrak{g}}_{1}^{\prime}}({\mathcal{O}}_{1})+\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}({\mathcal{O}}_{2})>\dim_{{\mathfrak{g}}_{1}^{\prime}}(\eta_{{\hat{\mathfrak{g}}_{1}^{\prime},\mathfrak{g}_{1}^{\prime}}}(\underline{p}(\psi^{1})))+\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}((\underline{p}(\psi_{2})^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}).$
@@
$\dim_{{\mathfrak{g}}_{1}^{\prime}}({\mathcal{O}}_{1})>\dim_{{\mathfrak{g}}_{1}^{\prime}}(\eta_{{\hat{\mathfrak{g}}_{1}^{\prime},\mathfrak{g}_{1}^{\prime}}}(\underline{p}(\psi^{1}))),$
@@
$\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}({\mathcal{O}}_{2})>\dim_{\hat{\mathfrak{g}}_{2}^{\prime}}((\underline{p}(\psi_{2})^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}).$
That is, either ${\mathcal{O}}_{1}$ is not related to or bigger than $\eta_{{\hat{\mathfrak{g}}_{1}^{\prime},\mathfrak{g}_{1}^{\prime}}}(\underline{p}(\psi^{1}))$, or, ${\mathcal{O}}_{2}$ is not related to or bigger than $(\underline{p}(\psi_{2})^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}$, which contradicts Conjecture 1.2 and Remark 1.4(3).
When ${\mathrm{G}}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}$, by Lemma 4.1, we can see that

### 06 May 2026 v2

@@ 5. Construction of elements in local Arthur packets
In the identity (6.1), we have that for any nilpotent orbit $\mathcal{O}$},
$c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})=\sum_{\pi\in\widetilde{\Pi}_{\psi}}c_{{\mathcal{O}}}(\pi).$
On the other hand, since $\underline{p}\in max\{\cup_{\pi\in\widetilde{\Pi}_{\psi}}\mathfrak{p}^{m}(\pi)\}$, by [MW87], for any nilpotent orbit ${\mathcal{O}}_{\underline{p}}$ corresponding to $\underline{p}$, $c_{{\mathcal{O}}_{\underline{p}}}(\pi)\in\mathbb{Z}_{\geq 0}$ and there exist $\pi\in\widetilde{\Pi}_{\psi}$ and a rational nilpotent orbit ${\mathcal{O}}_{0}\in{\mathcal{N}}_{\mathfrak{g}_{n}}$ corresponding to $\underline{p}$ such that $c_{{\mathcal{O}}_{0}}(\pi)\in\mathbb{Z}_{>0}$. Hence, $c_{{\mathcal{O}}_{0}}(\widetilde{\Pi}_{\psi})>0$. By Lemma 6.1, there exists a function $f_{0}\in\mathcal{H}(G_{n})$ with arbitrarily small support such that
@@
for any other ${\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}$. Let $f=f_{0}$ when $\mathrm{G}_{n}={\mathrm{Sp}}_{2n},{\mathrm{SO}}_{2n+1}$, and let $f=R(c_{{\mathrm{G}}_{n}})(f_{0}+f_{0}^{c})$ when $\mathrm{G}_{n}={\mathrm{SO}}_{2n}^{\alpha}$, where $c$ is the non-trivial element in the outer automorphism group. Then $f\in\widetilde{{\mathcal{H}}}(G_{n})$. When $\mathrm{G}_{n}={\mathrm{SO}}_{2n}^{\alpha}$, if ${\mathcal{O}}_{0}\neq{\mathcal{O}}_{0}^{c}$, i.e., ${\mathcal{O}}_{0}$ is a very even nilpotent orbit, then we have that
$\hat{\mu}_{{\mathcal{O}}_{0}}(f)=\hat{\mu}_{{\mathcal{O}}_{0}^{c}}(f)=1,\text{ and }\hat{\mu}_{{\mathcal{O}}}(f)=0,$
@@
Now, without loss of generality, we can assume that $f$, $R(c_{{\mathrm{G}}_{1}^{\prime}})f^{\prime 1}\otimes R(c_{{\mathrm{G}}_{2}^{\prime}})f^{\prime 2}$, $\widetilde{f}^{1}$, $\widetilde{f}^{2}$ , $\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}$, and $\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}$ all have small supports near 1. Take $t\in F^{\times}$ with $|t|$ small. By Lemma 6.2 and Lemma 6.3, we have
(6.2) $\begin{split}&\,\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{\mathcal{O}}((f)_{t^{2}})\\ =&\,|t|^{\dim(\mathfrak{g}_{n})-\dim(\mathfrak{g_{1}^{\prime}}^{\theta})-\dim(\mathfrak{g^{\prime}_{2}}^{\theta})}\left(\sum{{}_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g_{1}^{\prime}}^{\theta}}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}((\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}})_{t^{2}})\right)\\ &\,\cdot\left(\sum{{}_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g_{2}^{\prime}}^{\theta}}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}((\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}})_{t^{2}})\right).\end{split}$
@@
$\hat{\mu}_{\mathcal{O}}((f)_{t^{2}})$ $=|t|^{2\dim(\mathfrak{g_{n}})-\dim({\mathcal{O}})}\hat{\mu}_{\mathcal{O}}(f),$
$\hat{\mu}_{\mathcal{O}}((\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}})_{t^{2}})$ $=|t|{{}^{2\dim(\mathfrak{g_{1}^{\prime}}^{\theta})-\dim({\mathcal{O}})}}\hat{\mu}_{\mathcal{O}}(\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}),$
$\hat{\mu}_{\mathcal{O}}((\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}})_{t^{2}})$ $=|t|{{}^{2\dim(\mathfrak{g_{2}^{\prime}}^{\theta})-\dim({\mathcal{O}})}}\hat{\mu}_{\mathcal{O}}(\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}).$
Therefore, we obtain that
@@
(6.3) $\begin{split}&\,|t|^{\dim(\mathfrak{g}_{n})-\dim({\mathcal{O}}_{0})}c_{{\mathcal{O}}_{0}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{{\mathcal{O}}_{0}}(f)\\ =&\,\left(\sum{{}_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g_{1}^{\prime}}^{\theta}}}}|t|{{}^{\dim(\mathfrak{g_{1}^{\prime}}^{\theta})-\dim({\mathcal{O}})}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}(\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}})\right)\\ &\,\cdot\left(\sum{{}_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g_{2}^{\prime}}^{\theta}}}}|t|{{}^{\dim(\mathfrak{g_{2}^{\prime}}^{\theta})-\dim({\mathcal{O}})}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}(\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}})\right).\end{split}$
@@
Then one can see that, there exist nilpotent orbits ${\mathcal{O}}_{1}\in{\mathcal{N}}_{\mathfrak{g_{1}^{\prime}}^{\theta}}$, ${\mathcal{O}}_{2}\in{\mathcal{N}}_{\mathfrak{g_{2}^{\prime}}^{\theta}}$ such that $c_{{\mathcal{O}}_{1}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{{\mathcal{O}}_{1}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})$ and $c_{{\mathcal{O}}_{2}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{{\mathcal{O}}_{2}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})$ are all nonzero and
[…]
By Proposition 4.4, we can see that
[…]
@@
[…]
@@
[…]
contradicting Conjecture 1.2.
Next, we assume that for any $\pi\in\widetilde{\Pi}_{\psi}$, $\underline{p}\in\mathfrak{p}^{m}(\pi)$,
$\dim_{{\mathfrak{g}}_{n}}(\underline{p})\leq\dim_{{\mathfrak{g}}_{n}}(\eta_{{\hat{\mathfrak{g}}_{n},\mathfrak{g}_{n}}}(\underline{p}(\psi))).$
Assume that Conjecture 1.2 does not hold, i.e., there exists $\underline{p}\in\mathfrak{p}^{m}(\widetilde{\pi}_{\psi})$, such that
$\dim_{\mathfrak{g}_{n}^{\theta}}(\underline{p})>\dim_{\mathfrak{g}_{n}^{\theta}}((\underline{p}(\psi)^{t})_{{\mathrm{G}}_{n}^{\theta}}).$
Applying [Art13, Theorem 2.2.1, Part (b)] with $x=s=1$, we have the following distribution identity
(6.4) $),\sum_{\pi\in\widetilde{\Pi}_{\psi}}\langle s_{\psi},\pi\rangle f_{G_{n}}(\pi)=f(\psi)={\mathrm{tr}}(\widetilde{\pi}_{\psi}(\widetilde{f}$)),
where $f\in\widetilde{{\mathcal{H}}}(G_{n})$, $\widetilde{f}\in\widetilde{{\mathcal{H}}}(N)$ such that $\widetilde{f}$ transfers to $f^{G_{n}}$. We take the character expansions of both sides of (6.4). The left hand side is expanded at the identity, and the right hand side is expanded at the twisted points $\theta_{\widehat{\mathrm{G}}_{n}}$ (see (1.3)). We have the following equality:
(6.5) $\begin{split}\sum_{\pi\in\widetilde{\Pi}_{\psi}}\langle s_{\psi},\pi\rangle\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\pi)\hat{\mu}_{\mathcal{O}}(f)=\sum{{}_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}^{\theta}}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{n}}}}),\end{split}$
where $\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{n}}}$} is the Harish-Chandra descent of $\widetilde{f}$ (see for example [Kon02, Section 3.1]). Let $c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})=\sum_{\pi\in\widetilde{\Pi}_{\psi}}\langle s_{\psi},\pi\rangle c_{{\mathcal{O}}}(\pi)$, then (6.5) becomes
@@
(6.6) $\begin{split}\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{\mathcal{O}}(f)=\sum{{}_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}^{\theta}}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{n}}}}).\end{split}$
Since $\underline{p}\in\mathfrak{p}^{m}(\widetilde{\pi}_{\psi})$, there exists a rational nilpotent orbit ${\mathcal{O}}_{0}\in{\mathcal{N}}_{\mathfrak{g}_{n}^{\theta}}$ corresponding to $\underline{p}$ such that $c_{{\mathcal{O}}_{0}}(\widetilde{\pi})\neq 0$. Since both the transfer map and the Harish-Chandra descent map are surjective at 1 (see [Art13, Corollary 2.1.2] and [Var17, Proposition 3.9]), there exist $f$ and $\widetilde{f}$ such that $\hat{\mu}_{{\mathcal{O}}_{0}}({{}\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{n}}}})=1$ and $\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{n}}}})=0$ for any other ${\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}^{\theta}}$. Therefore, (6.6) becomes
$\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{\mathcal{O}}(f)=c_{{\mathcal{O}}_{0}}(\widetilde{\pi}_{\psi})\hat{\mu}_{{\mathcal{O}}_{0}}({{}\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{n}}}}).$
Now, without loss of generality, we can assume that $f$, $\widetilde{f}$, $\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{n}}}$ all have small supports near 1. Take $t\in F^{\times}$ with $|t|$ small. By Lemma 6.2 and Lemma 6.3, we have